# Single parameter aging and density scaling

### v1

@@ I Introduction
Physical aging has obvious implications for the performance and properties of glassy materials used in applications, such as precision molding [35, 36, 37, 38], oxide glasses used for optical fibers[39] and displays [40], plastics [41], polymers used for membranes[31, 42], bioactive glasses [43] etc., but aging also poses interesting questions for fundamental research. An intrinsic feature of aging is that it is non-linear and a central fundamental question is whether this non-linear out-of-equilibrium relaxation is governed by the same slow degrees of freedom that are responsible for the linear structural relaxation in the equilibrium state. This has been the prevailing conception [44, 45, 46, 27, 47, 48], but is possible that the physical aging mechanism could be governed by other processes. In particular, there are indications that for large up-jumps, aging proceeds in a heterogeneous way resembling a nucleation and growth rather than a gradual homogeneous softening [13, 7, 8]. Furthermore, there is evidence that for large down-jumps, a faster aging mechanism sets in at short aging times [49, 50, 2], while the long time aging (i.e., close to equilibrium) remains dominated by the slow structural relaxation[51].
@@
If (or when) the structural relaxation governs physical aging, then the out-of-equilibrium situation must be strongly connected to the the properties that govern the equilibrium structural relaxation. In general, there is no consensus as to what controls the structural relaxation time, but it has been established that many organic systems obey so-called density scaling [53, 54, 55, 56]. This means that the relaxation rate depends on a scaling parameter $\Gamma$, which in most cases is given by $\Gamma=\rho^{\gamma}/T$. For the systems where density scaling applies in equilibrium, it is natural to attempt to generalize the concept to out of equilibrium. One of us has earlier suggested how to do this [57, 58] and recently Di Lisio et al. [9] suggested an alternative generalization.
@@ II Material time and single parameter aging
In general, making predictions with the TNM formalism thus requires analytical expressions for the memory kernel, $M$, and for the out-of-equilibrium rate, $\tau(T,T_{f})$. Traditionally, non-exponential fitting functions have been proposed (such as the stretched exponential) for $M$ and various empirical expression for the out-of-equilibrium relaxation time, the most used being Moynihan’s [64]. However, even in equilibrium, we do not know the exact shape of $M$ or how $\tau$ behaves as a function of temperature (in fact these are the two big questions in glass science), so this introduces as a minimum 4 different fitting parameters depending on the choice of fitting functions.
Single parameter aging (SPA)[59] is a strategy to avoid fitting functions in the description of temperature jump experiments, but in its essence not different from the TNM formalism. The fundamental idea of SPA is that all properties of the aging material are governed by a single non-equilibrium parameter. It follows that there is a one-to-one connection between any two aging properties, including the instantaneous relaxation time itself and some measured property, $X$. The simplest assumption for this relation is linearity, i.e., that logarithm of the aging rate and the measured property are linearly connected, $\ln\tau(t)-\ln\tau_{\text{eq}}=a(X(t)-X_{\text{eq}})=a\Delta X(t)$. Assuming furthermore that the total change in the measured property is linear with changes in temperature, $\Delta X(0)=b\Delta T$, i.e., the difference between the equilibrium value at the initial temperature before the jump and the equilibrium value at the new temperature, lead to the following expression for the aging rate
@@
If SPA applies, then it is possible to predict the aging curves of any temperature jump from just one aging curve and the fitting parameter $c$ if the final equilibrium relaxation times, $\tau_{\text{eq}}$, are known. We make use of this to test SPA. The principle is outlined in the following and described in full detail in the Supplementary Material.
The consequence of SPA (and also of the TNM formalism) is that the relaxation curves from all jumps collapse when plotted as a function material time, $R(\xi)$. Obviously, this can also be reversed such that the material time of all jumps are identical functions of the relaxation function $\xi(R)$. Knowing the material time and the out-of-equilibrium relaxation time, we can obtain the laboratory time corresponding to a given temperature jump by isolating and integrating Eq. 2.
We have
$d\xi=\frac{dt_{1}}{\tau_{1}}=\frac{dt_{2}}{\tau_{2}}\quad\Rightarrow\quad dt_{2}=\tau_{2}\,d\xi=\frac{\tau_{2}}{\tau_{1}}dt_{1}\,,$ (5)
where the subscripts 1 and 2 refer to two different temperature jumps, 1 and 2. Inserting Eq. (3) in Eq. 5 and integrating we obtain the time axis for temperature jump 2
@@
Thus, given $c$ and the equilibrium relaxation times, $\tau_{\text{eq},n}$, one measured temperature-jump relaxation curve can predict all other by transforming the time axis.
@@ III SPA analysis of temperature jumps
Di Lisio et al.[9] measured a series of temperature down-jumps of sizes up to 22.5 K for five different organic liquids: PDE, KDE, OTP, BMPC, and BMMPC using Fast Scanning Calorimetry. They used the enthalpy overshoot in heating to extract the fictive temperature, $T_{f}$, after a different waiting times at the annealing temperature, $T_{a}$. In this way they constructed curves for the temporal evolution of $T_{f}$ during aging. Subsequently, these curves were analyzed in the SPA framework combined with a density scaling assumption.
It is by no means surprising if SPA did not work for the large temperature down-jumps presented by Di Lisio et al. However, applying the SPA in this simplest form, i.e. with linearity assumptions and $c$ as a sole fitting parameter for all curves, we demonstrate below that SPA does capture all of the measured data, even the largest down-jumps reported by Di Lisio et al.
@@
Figure 1(A-E) shows the resulting SPA analysis (see Supplemental Material for details of the analysis). For each liquid, one jump was selected as the predictor of the rest. Since the data are sparse, we fitted the selected jump with a stretched exponential to serve as interpolation between data points. The curve fitted to the chosen jump is shown as a dashed line. This curve is then used to generate the other temperature jumps from Eq. (6) and the results are shown as full lines. For all data sets and all temperatures, there is a decent agreement between the predicted curve and and the measured data points. There are small deviations from the predictions for the largest temperature jumps for KDE, BMMPC, and BMPC at long times. For BMPC the predicted curve for the largest temperature jump is consistently below the data points. However, this is a simple scaling and can be explained by uncertainty in determining the equilibrium relaxation time from the data (see Supplementary Material for more details). Overall, SPA works to a remarkably good degree, even for the largest temperature jumps of $\Delta T>20$ K, when conducted in its original and simplest form. Figure 1(f) shows $g(R)$ as a function of $R$ for the largest jumps for each liquid. The simple SPA $g(R)$ is linear by definition.
@@ IV Combining single parameter aging with density scaling
To get an expression for the relaxation time during aging in this framework is equilibrium density scaling in the original version of Alba-Simionesco, Tarjus, and Kivelson [53, 54], where $\Gamma$ is defined as $\Gamma=e(\rho)/T$ with $e(\rho)$ interpreted as an energy scale that only depends on density while the topology of the energy landscape depends on both temperature and density though $\Gamma$. This leads to the following expression for the activation energy in equilibrium:
@@
When generalizing to out of equilibrium we assume that the energy scale $e(\rho)$ is controlled by density while the topology of the energy landscape is controlled by the non-equilibrium structural state of the system, described by the paramter, $\Gamma_{fic}$. This parameter can be expressed as $\Gamma_{fic}=e(\rho)/T_{\Gamma f}$, where $e(\rho)$ is defined as in equilibrium and $T_{\Gamma f}$ is the $\Gamma_{fic}$-temperature (Gamma-fictive temperature), which is conceptually similar to the classical fictive temperature, though it has a quite different behavior [58].
@@
The ansatz in this framework is that $\Gamma_{fic}$ which is connected to the microscopic structure, needs time to change. Right after a temperature jump it will therefore be equal to the value it had before the jump was performed. Thus, Eq. (11) predicts that the relaxation time is further from equilibrium at short times than predicted by Di Lisio et al.. In the case of down-jumps in temperature it means that the curves predicted by Eq. (11) will be more stretched than those predicted by Di Lisio et al. – and it is in fact seen in Fig. 2 that the data are more stretched than predicted by Di Lisio et al. Hénot et al. [16] have earlier this year made a similar qualitative comparison and likewise find that Eq. (11) has a better chance of representing the data. Hénot et al. consider up-jumps, and the stronger history dependence in this case means that Eq. (11) predicts longer relaxation times right after the temperature jump is performed, which will lead to a more compressed curve.
@@ Acknowledgements
We thank Jeppe C. Dyre and Nicholas P. Bailey for discussions of the manuscript and Valerio Di Lisio and Daniele Cangialosi for kindly sharing their data. This work was supported by the VILLUM Foundation’s Matter Grant No. VIL16515.

### v2

@@ I Introduction
Physical aging has obvious implications for the performance and properties of glassy materials used in applications, such as precision molding [35, 36, 37, 38], oxide glasses used for optical fibers[39] and displays [40], plastics [41], polymers used for membranes[31, 42], bioactive glasses [43] etc., but aging also poses interesting questions for fundamental research. An intrinsic feature of aging is that it is non-linear and a central fundamental question is whether this non-linear out-of-equilibrium relaxation is governed by the same slow degrees of freedom that are responsible for the linear structural relaxation in the equilibrium state. This has been the prevailing conception [44, 45, 46, 27, 47, 48], but it is possible that the physical aging mechanism could be governed by other processes. In particular, there are indications that for large up-jumps, aging proceeds in a heterogeneous way resembling a nucleation and growth rather than a gradual homogeneous softening [13, 7, 8]. Furthermore, there is evidence that for large down-jumps, a faster aging mechanism sets in at short aging times [49, 50, 2], while the long time aging (i.e., close to equilibrium) remains dominated by the slow structural relaxation[51].
@@
If (or when) the structural relaxation governs physical aging, then the out-of-equilibrium situation must be strongly connected to the properties that govern the equilibrium structural relaxation. In general, there is no consensus as to what controls the structural relaxation time, but it has been established that many organic systems obey so-called density scaling [53, 54, 55, 56]. This means that the relaxation rate depends on a scaling parameter $\Gamma$, which in most cases is given by $\Gamma=\rho^{\gamma}/T$. For the systems where density scaling applies in equilibrium, it is natural to attempt to generalize the concept to out of equilibrium. One of us has earlier suggested how to do this [57, 58] and recently Di Lisio et al. [9] suggested an alternative generalization.
@@ II Material time and single parameter aging
In general, making predictions with the TNM formalism thus requires analytical expressions for the memory kernel, $M$, and for the out-of-equilibrium relaxation time, $\tau(T,T_{f})$. Traditionally, non-exponential fitting functions have been proposed (such as the stretched exponential) for $M$ and various empirical expression for the out-of-equilibrium relaxation time, the most used being Moynihan’s [64]. However, even in equilibrium, we do not know the exact shape of $M$ or how $\tau$ behaves as a function of temperature (in fact these are the two big questions in glass science), so this introduces as a minimum 4 different fitting parameters depending on the choice of fitting functions.
Single parameter aging (SPA)[59] is a strategy to avoid fitting functions in the description of temperature jump experiments, but in its essence not different from the TNM formalism. The fundamental idea of SPA is that all properties of the aging material are governed by a single non-equilibrium parameter. It follows that there is a one-to-one connection between any two aging properties, including the instantaneous relaxation time itself and some measured property, $X$. The simplest assumption for this relation is linearity, i.e., that logarithm of the out-of-equilibrium relaxation time and the measured property are linearly connected, $\ln\tau(t)-\ln\tau_{\text{eq}}=a(X(t)-X_{\text{eq}})=a\Delta X(t)$. Assuming furthermore that the total change in the measured property is linear with changes in temperature, $\Delta X(0)=b\Delta T$, i.e., the difference between the equilibrium value at the initial temperature before the jump and the equilibrium value at the new temperature, lead to the following expression for the out-of-equilibrium relaxation time
@@
If TNM formalism applies, the relaxation curves from all jumps collapse when plotted as a function material time by a transformation of the time axis according to Eq. (2). The transformation can also be reversed such that – if the material time and the out-of-equilibrium relaxation time $\tau(t)$ corresponding to a given temperature jump is known – we can obtain the laboratory time corresponding to that jump. In this sense, one can think of $R$ as the independent variable and time as the dependent variable. This is true for any temperature jump and consequently we can transform the time axis from one non-linear jump to another. In a condensed form, the above can be written as
$\begin{split}\frac{dt_{1}(R)}{\tau_{1}}&=d\xi(R)=\frac{dt_{2}(R)}{\tau_{2}}\Rightarrow\\ dt_{2}&(R)=\frac{\tau_{2}}{\tau_{1}}dt_{1}(R)\,,\end{split}$ (5)
where the subscripts 1 and 2 refer to two different temperature jumps, 1 and 2.
SPA gives a relation between the out-of-equilibrium relaxation time and the measured quantity (Eq. (3)), which inserted in Eq. (5) leads to the time axis for temperature jump 2
@@
Thus, given $c$ and the equilibrium relaxation times, $\tau_{\text{eq},n}$, the SPA provides a recipe for obtaining any temperature-jump relaxation curve from a given measured temperature jump.
@@ III SPA analysis of temperature jumps
Di Lisio et al. measured a series of temperature down-jumps of sizes up to 22.5 K using fast scanning calorimetry for five different organic liquids: phenolphthalein dimethyl ether (PDE), o-Cresolphthalein dimethyl ether (KDE), o-terphenyl (OTP), bisphenol-C-dimethylether (BMPC), and 1,1-bis (4-methoxyphenyl)cyclohexane (BMMPC). They used the enthalpy overshoot in heating to extract the fictive temperature, $T_{f}$, after a different waiting times at the annealing temperature, $T_{a}$. In this way they constructed curves for the temporal evolution of $T_{f}$ during aging. Subsequently, these curves were analyzed in the SPA framework combined with a density scaling assumption.
It is by no means surprising if SPA did not work for the large temperature down-jumps presented by Di Lisio et al. However, applying SPA in this simplest form, i.e. with linearity assumptions and $c$ as a sole fitting parameter for all curves, we demonstrate below that SPA does capture all of the measured data, even the largest down-jumps reported by Di Lisio et al.
@@
Figure 1(A-E) shows the resulting SPA analysis (see Supplemental Material for details of the analysis). For each liquid, one jump was selected as the predictor of the rest. Since the data are sparse, we fitted the selected jump with a stretched exponential to serve as interpolation between data points. The curve fitted to the chosen jump is shown as a dashed line. This curve is then used to generate the other temperature jumps from Eq. (6) and the results are shown as full lines. For all data sets and all temperatures, there is a decent agreement between the predicted curve and and the measured data points. The fitted values for $c$ is given in Table 1.
Liquid $c$ (K-1)
KDE 0.30
PDE 0.30
OTP 0.55
BMMPC 0.37
BMPC 0.46
Table 1: Fitted values for $c$ of Eq. (6).
There are small deviations from the predictions for the largest temperature jumps for KDE, BMMPC, and BMPC at long times. For BMPC the predicted curve for the largest temperature jump is consistently below the data points. However, this is a simple scaling and can be explained by uncertainty in determining the equilibrium relaxation time from the data (see Supplementary Material for more details). Overall, SPA works to a remarkably good degree, even for the largest temperature jumps of $\Delta T>20$ K, when conducted in its original and simplest form. Figure 1(f) shows $g(R)$ as a function of $R$ for the largest jumps for each liquid. The simple SPA $g(R)$ is linear by definition.
@@ IV Combining single parameter aging with density scaling
To get an expression for the relaxation time during aging in this framework, the starting point is the equilibrium density scaling in the original version of Alba-Simionesco, Tarjus, and Kivelson [53, 54], where $\Gamma$ is defined as $\Gamma=e(\rho)/T$ with $e(\rho)$ interpreted as an energy scale that only depends on density while the topology of the energy landscape depends on both temperature and density through $\Gamma$. This leads to the following expression for the activation energy in equilibrium:
@@
When generalizing to out-of-equilibrium situation we assume that the energy scale $e(\rho)$ is controlled by density while the topology of the energy landscape is controlled by the non-equilibrium structural state of the system, described by the paramter, $\Gamma_{fic}$. This parameter can be expressed as $\Gamma_{fic}=e(\rho)/T_{\Gamma f}$, where $e(\rho)$ is defined as in equilibrium and $T_{\Gamma f}$ is the $\Gamma_{fic}$-temperature (Gamma-fictive temperature), which is conceptually similar to the classical fictive temperature, though it has a quite different behavior [58].
@@
The ansatz in this framework is that $\Gamma_{fic}$ which is connected to the microscopic structure, needs time to change. Right after a temperature jump it will therefore be equal to the value it had before the jump was performed. Thus, Eq. (11) predicts that the relaxation time is further from equilibrium at short times than predicted by Di Lisio et al.. In the case of down-jumps in temperature it means that the curves predicted by Eq. (11) will be more stretched than those predicted by Di Lisio et al. consistent with the data in Fig. 2 being more stretched than the prediction. Hénot et al. [16] have earlier this year made a similar qualitative comparison and likewise find that Eq. (11) has a better chance of representing the data. Hénot et al. consider up-jumps, and in this case it means that Eq. (11) predicts longer relaxation times right after the temperature jump is performed than predicted by the Di Lisio expression, leading to a more compressed curve.
@@ Acknowledgements
We thank Jeppe C. Dyre and Nicholas P. Bailey for discussions of the manuscript. Furthermore, we are grateful for the continued open discussions with Daniele Cangialosi about aging in general and SPA in particular. Finally, we would like to thank Valerio Di Lisio and Daniele Cangialosi for generously sharing their data. This work was supported by the VILLUM Foundation’s Matter Grant No. VIL16515.